If the police have 8 suspects, how many different ways can they select 5 for a lineup?
The answer that I came up with was 6720 ways. But its wrong. Was i supposed to do this problem as a permutation or a combination? Or neither?

Answer by stanbon(75887)

If the police have 8 suspects, how many different ways can they select 5 for a lineup?
The answer that I came up with was 6720 ways. But its wrong. Was i supposed to do this problem as a permutation or a combination? Or neither?
-----
You are selecting groups so use a combination formula.
8C5 = 8C3 = (8*7*6)/(1*2*3) = 8*7 = 56
---------
Cheers,
Stan H.
------------------------